# Revisiting Gamma-Ray Emission of the Supernova Remnant RCW 103

### v1

@@ Abstract
We analyze more than 15 years of the -ray data, obtained with the Large Area Telescope (LAT) onboard the Fermi Gamma-ray Space Telescope (Fermi), for the region of the young supernova remnant (SNR) RCW 103, since the nearby source 4FGL J1616.25054e, counterpart to HESS J1616518 and 13 arcmin away from the SNR, is determined to be extended in the more recent Fermi-LAT source catalog. Different templates for 4FGL J1616.25054e and RCW 103 are tested, and we find that a point source with a power-law (PL) spectrum at the southern limb of the SNR best describes the detected -ray emission. The photon index of the PL emission is , softer than the previously reported when the counterpart to HESS J1616518 was considered to be a point source (which likely caused mis-identification of extended emission at RCW 103). In order to produce the -ray emission in a hadronic scenario, we estimate that protons with an index2.4 PL energy distribution are needed. These results fit with those from multi-wavelength observations that have indicated the remnant at the southern limb is interacting with a molecular cloud.
@@ 2.2 Maximum Likelihood Analysis
helps reveal the emission in a high, spatial resolution222http://www.slac.stanford.edu/exp/glast/groups/canda
/lat_Perfomance.htm
@@ 2.3 Spatial Distribution Analysis and Source Identification
a point source (PS; CCO model), a uniform disk with a radius of
@@
analysis to the data with each of the models, CCO HESS0.32,
@@
To be as completely as possible, we re-checked the extension of the counterpart of HESS J1616508 by setting Radiops for RCW 103. The uniform disk with a radius of 002–050 (in a 002 step) was tested again. We found that a 038 radius (thus the HESS0.38 model used above) resulted in a maximum value and the fit improvement of the model of Radiops HESS0.38 is at a 4.9 significance level with respect to that of Radiops HESS0.32 (Table 1). We thus considered this Radiops HESS0.38 model as the final one and used it in the following analysis.
@@ 2.4 Spectral Analysis
The obtained spectra with both statistical and systematic uncertainties are shown in Figure 3, while the flux and TS values of the spectral data points are provided in Table 2. The PL spectral models obtained for the two sources from the likelihood analysis match the spectral data points.
@@ 3 Discussion
Using the latest Fermi-LAT source catalog, we have revisited the detection of -ray emission of RCW 103 by analyzing the data collected for years. The SNR’s emission is resolved at high-energy bands (Figures 1 & 2), even though the nearby source 4FGL J1616.25054e, the counterpart to the VHE source HESS J1616508, has been determined to be extended. Different from the previous result reported in Xing et al. (2014) that the emission was extended, we have found that it is consistent with being a PS. It has a PL spectrum with photon index , and this PL is softer than the previously reported ( in Xing et al. 2014). We note that this spectrum is very similar to that of the SNR Kes 17, which had (Supán et al., 2023).

### v2

@@ Abstract
We analyze more than 15 years of the -ray data, obtained with the Large Area Telescope (LAT) onboard the Fermi Gamma-ray Space Telescope (Fermi), for the region of the young supernova remnant (SNR) RCW 103, since the nearby source 4FGL J1616.25054e, counterpart to HESS J1616518 and 13 arcmin away from the SNR, is determined to be extended in the more recent Fermi-LAT source catalog. Different templates for 4FGL J1616.25054e and RCW 103 are tested, and we find that a point source with a power-law (PL) spectrum at the southern limb of the SNR best describes the detected -ray emission. The photon index of the PL emission is (or in a Log-Parabola model), softer than the previously reported when the counterpart to HESS J1616518 was considered to be a point source (which likely caused mis-identification of extended emission at RCW 103). In order to produce the -ray emission in a hadronic scenario, we estimate that protons with an index2.4 PL energy distribution are needed. These results fit with those from multi-wavelength observations that have indicated the remnant at the southern limb is interacting with a molecular cloud.
@@ 2.2 Maximum Likelihood Analysis
helps reveal the emission in a high, spatial resolution222https://www.slac.stanford.edu/exp/glast/groups/canda
/lat_Performance.htm
@@ 2.3 Spatial Distribution Analysis and Source Identification
a point source (PS; PS model), a uniform disk with a radius of
@@
analysis to the data with each of the models, PS HESS0.32,
@@
To be as complete as possible, we re-checked the extension of the counterpart of HESS J1616508 by setting Radiops for RCW 103. The uniform disk with a radius of 002–050 (in a 002 step) was tested again. We found that a 038 radius (thus the HESS0.38 model used above) resulted in a maximum value and the fit improvement of the model of Radiops HESS0.38 is at a 4.9 significance level with respect to that of Radiops HESS0.32 (Table 1). We thus considered this Radiops HESS0.38 model as the final one and used it in the following analysis.
@@
We also tested a Log-Parabola model, , for the emission of RCW 103. The parameters and were set free, and , a scale parameter, was always fixed. Different values were tested. If we fixed it at 5 GeV, the resulting likelihood value indicated the fitting was better than that with the PL at a 2 significance level, where , , and photon cm-2 s-1 (TS ). The results are similar to those with the PL (note that the curvature parameter is very small), but the uncertainties are larger.
@@ 2.4 Spectral Analysis
The obtained spectra with both statistical and systematic uncertainties are shown in Figure 3, while the flux and TS values of the spectral data points are provided in Table 2. The PL spectral models obtained for the two sources from the likelihood analysis match the spectral data points. However for RCW 103, the two upper limits in 17–91 GeV (Table 2) are at the model-fit line, showing a slight inconsistency. Thus, its emission may be better described with the Log-Parabola model we obtained in Section 2.3, which is also shown in Figure 3.
@@ 3 Discussion
Using the latest Fermi-LAT source catalog, we have revisited the detection of -ray emission of RCW 103 by analyzing the data collected for years. The SNR’s emission is resolved at high-energy bands (Figures 1 & 2), even though the nearby source 4FGL J1616.25054e, the counterpart to the VHE source HESS J1616508, has been determined to be extended. Differing from the previous result reported in Xing et al. (2014) that the emission was extended, we have found that it is consistent with being a PS. It has a PL spectrum with photon index or a Log-Parabola one with slope , and the spectrum is softer than the previously reported ( in Xing et al. 2014). We note that this spectrum is very similar to that of the SNR Kes 17, which had (Supán et al., 2023).